OpenStudy (anonymous):

Simplify the expression 2003-02-03-00-00_files/i0080000.jpg.

OpenStudy (anonymous):

lol alright.

OpenStudy (anonymous):

no not that lol

OpenStudy (anonymous):

I know, I was kidding.

OpenStudy (anonymous):

let me type the problem

OpenStudy (anonymous):

\[\frac{ ^{2^{5}\times 3^{4}} }{ 2^{7}\times 3^{3}? }\]

OpenStudy (anonymous):

u get that?

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

ya sorry, was answering someone elses question

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

so 2^5 =32 and 3^4 =81

OpenStudy (anonymous):

32 x 81 = 2592 in the numerator.

OpenStudy (anonymous):

2^7 = 128 AND 3^3=27

OpenStudy (anonymous):

so 128 x 27 =3456 in denominator

OpenStudy (anonymous):

i got to simplify. how

OpenStudy (anonymous):

so 2592/3456 =.75 or 3/4

OpenStudy (anonymous):

you just plug the numbers into a calculator to simplify (always remember PEMDAS)

OpenStudy (anonymous):

the answers are fractions

OpenStudy (anonymous):

so it's 3/4

OpenStudy (anonymous):

thats not 1 of the answers

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

It's 100% correct, I know.